Answer the following:   

a. byte a = 10, b = 6;

byte c = a ^ c;

What is the value of a?

b. byte a = 12, b = 6;

byte c = a ^ b;

     What is the value of c?
Answer #1

A XOR B 1

Part a.)
(Assuming that the operation is not c = c^a as per the problem, as this will give an error as the initial value of c has not been defined)

Binary Representation
Value of a = 1 0 1 0
Value of b = 0 1 1 0
--------------------------
Value of c = 1 1 0 0
[c = a^b (^ - Bitwise XOR operation)]

Therefore as per XOR table shown above c = 1100

Hence value of c is 12 in PART a.)


Part b.)

Binary Representation
Value of a = 1 1 0 0
Value of b = 0 1 1 0
--------------------------
Value of c = 1 0 1 0
[c = a^b (^ - Bitwise XOR operation)]

Therefore as per XOR table shown above c = 1010

Hence value of c is 10 in PART b.)